Problem 3-11 T-Account Analysis of Cost Flows [LO3-2, LO3-3, LO3-4]

Selected T-accounts of Moore Company are given below for the just completed year:

Raw Materials
Bal. 1/1 17,000 Credits ?
Debits 124,000
Bal. 12/31 27,000
Manufacturing Overhead
Debits 229,800 Credits ?
Work in Process
Bal. 1/1 22,000 Credits 476,000
Direct materials 92,000
Direct labor 156,000
Overhead 241,800
Bal. 12/31 ?
Factory Wages Payable
Debits 189,000 Bal. 1/1 10,000
Credits 184,000
Bal. 12/31 5,000
Finished Goods
Bal. 1/1 42,000 Credits ?
Debits ?
Bal. 12/31 63,000
Cost of Goods Sold
Debits ?

Required:

1. What was the cost of raw materials used in production during the year?

2. How much of the materials in (1) above consisted of indirect materials?

3. How much of the factory labor cost for the year consisted of indirect labor?

4. What was the cost of goods manufactured for the year?

5. What was the unadjusted cost of goods sold for the year? Do not include any underapplied or overapplied overhead in your answer.

6. If overhead is applied to production on the basis of direct labor cost, what predetermined overhead rate was in effect during the year?

7. Was manufacturing overhead underapplied or overapplied? By how much?

8. Compute the ending balance in Work in Process. Assume that this balance consists entirely of goods started during the year. If $8,500 of this balance is direct labor cost, how much of it is direct materials cost? Applied overhead cost?
Answer #1

1) Cost of raw material used = 17000+124000-27000 = 114000

2) Indirect material = 114000-92000 = 22000

3) Indirect labor = 184000-156000 = 28000

4) Cost of goods manufactured = 476000

5) Unadjusted cost of goods sold = 42000+476000-63000 = 455000

6) Overhead rate = 241800/156000 = 155% of labor cost

7) Applied overhead = 241800

Actual overhead = 229800

Over applied overhead = 241800-229800 = 12000

8)Ending work in process = 35800

Applied overhead = 8500*1.55 = 13175

Direct material = 35800-8500-13175 = 14125